OpenStudy (anonymous):

Best Response Given.!!!! 7(t + 8.5) - 5t +4 *Simplify The Expression* Also: 4.3(5.6 + c) +9c 16b - 4(c + 3) - 4b

OpenStudy (anonymous):

I Would Fan But My Laptop's Messed Up. :c

OpenStudy (kc_kennylau):

I'll give you an example which can help you to do the rest :) E.g.: 14e+3.4(e+2)-3e \[14e+3.4(e+2)-3e\]\[=14e+3.4e+3.4\times2-3e\]\[=14e+3.4e-3e+6.8\]\[=14.4e+6.8\] Learn more: https://www.youtube.com/watch?v=YZBStgZGyDY